OpenStudy (anonymous):

Find the ratio in simplest form. 30 minutes to 2 hours? A. 1/4 B. 1/15 C. 15/1

OpenStudy (anonymous):

2 hours= 2*60 minutes= 120 minutes so, 30 minutes to 2 hours = 30 /120 = 1/4=1:4

OpenStudy (maheshmeghwal9):

1 hr. = 60min. 2hr. = 120 min. 30min/2ghrs=30min/120min. =>1/4

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (maheshmeghwal9):

yw :) but in my solution A typing mistake {2hrs*}